All Systems Operational
Regions Operational
Global   Operational
AMS2   Operational
AMS3   Operational
BLR1   Operational
FRA1   Operational
LON1   Operational
NYC1   Operational
NYC2   Operational
NYC3   Operational
SFO1   Operational
SFO2   Operational
SGP1   Operational
TOR1   Operational
Services Operational
API   Operational
Block Storage   Operational
Cloud Control Panel   Operational
Cloud Firewall   Operational
Community   Operational
DNS   Operational
Droplets   Operational
Event Processing   Operational
Load Balancers   Operational
Monitoring   Operational
Networking   Operational
Spaces   Operational
Support Center   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Aug 15, 2018

No incidents reported today.

Aug 14, 2018
Resolved - Our engineering team has resolved the issue with Block Storage in our NYC3 region. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 14, 15:03 UTC
Monitoring - Our engineering team has implemented a fix to the issue with Block Storage in our NYC3 region and is monitoring the results. We will post an update as soon as the issue is fully resolved.
Aug 14, 14:33 UTC
Identified - Our engineering team has identified the cause of the issue with Block Storage in our NYC3 region and is actively working on a fix. We will post an update as soon as additional information is available.
Aug 14, 14:15 UTC
Investigating - Our engineering team is investigating an issue with Block Storage in our NYC3 region. During this time you may experience issues with accessing and managing Block Storage. We apologize for the inconvenience and will share an update once we have more information.
Aug 14, 13:43 UTC
Aug 13, 2018

No incidents reported.

Aug 12, 2018

No incidents reported.

Aug 11, 2018

No incidents reported.

Aug 10, 2018

No incidents reported.

Aug 9, 2018
Resolved - Our engineering team has resolved the issue with networking in our NYC3 region. If you continue to experience any issues with packet loss or dropped connections, please feel free to reach out to us through the Support Portal. We apologize for the inconvenience.
Aug 9, 21:00 UTC
Monitoring - Networking issues are no longer impacting our NYC3 region. Users should no longer be experiencing packet loss or dropped connections. We are monitoring the situation and will post an update as soon as the issue is fully resolved.
Aug 9, 20:12 UTC
Investigating - Our engineering team is investigating a networking issue in our NYC3 region. At this time, you may experience packet loss or dropped connections. We apologize for the inconvenience and will share an update once we have more information.
Aug 9, 19:22 UTC
Resolved - Our engineering team has resolved the issue with emails not being received by our Support team. If you continue to experience problems with emailing support, please feel free to reach out to us through the Support Portal. We apologize for the inconvenience and will publish a postmortem to provide more detail about this issues once we've completed our internal review.
Aug 9, 20:25 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with emails sent to our Support team. Emails sent during the course of the incident, and any new emails being sent, are now being processed, though there may be a delay as our system processes the backlog of emails. We are continuing to monitor the situation and we will post an update as soon as we confirm full resolution.
Aug 9, 19:50 UTC
Investigating - Our engineering team is investigating an issue with emails sent to our Support team not being properly received. If you have emailed our Support team after 8:32PM UTC yesterday, we may not have received your email. The alternate method of reaching our Support team is through the Support Portal; however, Salesforce is experiencing an unrelated issue causing significant delays with tickets being submitted through the Support Portal. We apologize for the inconvenience and will share an update once we have more information.
Aug 9, 19:09 UTC
Resolved - Our engineering team has resolved the issue with Floating IPs. The process of unassigning these Floating IPs should now be operating normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 9, 14:16 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue impacting Floating IPs not being unassigned upon request and we are monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 9, 13:33 UTC
Update - Our engineering team continues to investigate the issue with Floating IPs not unassigning in the Cloud Control Panel until a refresh is performed. Thank you for your patience. We will post an update as soon as additional information is available.
Aug 9, 12:33 UTC
Update - Our engineering team continues to investigate the issue with Floating IPs. We appreciate your patience and will post an update as soon as additional information is available.
Aug 9, 11:51 UTC
Investigating - Our engineering team is investigating an issue impacting Floating IPs. During this time, you may experience delays in un-assigning Floating IPs on Droplets. We apologize for the inconvenience and will share an update once we have more information.
Aug 9, 10:58 UTC
Completed - The scheduled maintenance has been completed.
Aug 9, 05:10 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 9, 02:10 UTC
Scheduled - Our networking team will be upgrading core networking equipment to improve performance and reliability for our TOR1 location.

There will be a few minutes of event delays in the datacenter while hardware is cutover. No impact of internet bound traffic is expected.

If you have any questions or concerns, please reach out to our support team by opening up a ticket on your account.
Aug 9, 02:08 UTC
Aug 8, 2018
Completed - The scheduled maintenance has been completed. If you have any questions or concerns, please reach out to our support team by opening up a ticket on your account
Aug 8, 18:12 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 8, 17:01 UTC
Scheduled - Our networking team will be upgrading core networking equipment to improve performance and reliability for our BLR1 location.

There will be a few minutes of event delays in the datacenter while hardware is cutover. No impact of internet bound traffic is expected.

If you have any questions or concerns, please reach out to our support team by opening up a ticket on your account.
Aug 8, 16:30 UTC
Resolved - Our engineering team has resolved the issue with delayed Droplet destroy event processing. Destroy events should now be processing normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 8, 14:53 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with delayed Droplet destroy event processing and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 8, 14:23 UTC
Identified - Our engineering team has identified the cause of the issue with delayed Droplet destroy events and is actively working on a fix. We will post an update as soon as additional information is available.
Aug 8, 13:45 UTC
Aug 7, 2018
Completed - The scheduled maintenance has been completed.
Aug 7, 23:26 UTC
Update - Scheduled maintenance is still in progress. We will provide updates as necessary.
Aug 7, 23:26 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 7, 21:00 UTC
Scheduled - Our Spaces engineering team will be performing maintenance on elements of the Spaces components in NYC3.

Customers may experience intermittent, short-lived drops in throughput, though no significant performance impact is expected. Updates will follow as work progresses.
Aug 7, 20:43 UTC
Aug 6, 2018

No incidents reported.

Aug 5, 2018

No incidents reported.

Aug 4, 2018
Resolved - Our engineering team has resolved the issue with networking in our BLR1 region. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 4, 23:07 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with networking in our BLR1 region and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 4, 22:42 UTC
Identified - Our engineering team has identified the cause of the current networking issues in the BLR1 region. At this time all events in the BLR1 region are impacted including droplet creation events.
Aug 4, 21:48 UTC
Investigating - Our engineering team is investigating an issue with networking in our BLR1 region. Over the past 15 minutes you may experience packet loss or dropped connections. We apologize for the inconvenience and will share an update once we have more information.
Aug 4, 21:30 UTC
Aug 3, 2018
Resolved - Our engineering team has resolved the issue with event processing. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 3, 15:47 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with event processing delays and is monitoring the situation. Events should now be processing normally. We will post an update as soon as the issue is fully resolved.
Aug 3, 15:07 UTC
Update - Our engineering team continues to investigate the issue with delayed event processing. The scope of impacted events has broadened since our last update; at this time, users may experience delays when attempting to create or destroy Droplets, and other events may experience delays as well. We apologize for the inconvenience and will share an update once we have more information.
Aug 3, 14:32 UTC
Investigating - Our engineering team is investigating an issue with delayed event processing. During this time, you may experience Droplet destroy actions taking longer than usual. We apologize for the inconvenience and will share an update once we have more information.
Aug 3, 14:12 UTC
Resolved - Our engineering team has resolved the issue with Metadata Processing. Metadata Processing should now be operating normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 3, 11:51 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with Metadata Processing and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 3, 11:17 UTC
Identified - Our engineering team has identified the cause of the issue with Metadata Processing and is actively working on a fix. We will post an update as soon as additional information is available.
Aug 3, 10:53 UTC
Investigating - Our engineering team is investigating an issue with delays in Metadata Processing. We apologize for the inconvenience and will share an update once we have more information.
Aug 3, 10:23 UTC
Aug 2, 2018
Postmortem - Read details
Aug 7, 17:50 UTC
Resolved - Our engineering team has resolved the issue with accessing the Billing area. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 2, 13:38 UTC
Monitoring - Our engineering team has confirmed that the fix implemented last night has mitigated errors received when users attempted to access the Billing area. We are monitoring the situation and we will post an update as soon as the issue is fully resolved.
Aug 2, 12:55 UTC
Update - Our engineering team has implemented a fix to stabilize the intermittent errors some users experienced when attempting to access the Billing area in the Cloud Control Panel. We are continuing to monitor the situation throughout the night, and will post an update in the morning (ET) once we can confirm full resolution.
Aug 2, 01:37 UTC
Update - Our engineering team is working to resolve the issue causing intermittent errors for some users attempting to access the Billing area in the Cloud Control Panel. We appreciate your patience and will share additional updates as they become available.
Aug 2, 00:57 UTC
Identified - While our engineering team has implemented a fix to resolve the issue causing errors on the Cloud Control Panel, we have identified a related issue causing errors for some users attempting to access or view the Billing area. We are working towards a solution and will share additional updates as they become available.
Aug 1, 22:54 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issues affecting Cloud Control Panel, and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 1, 21:58 UTC
Identified - Our engineering team has identified the issue causing errors within Cloud Control Panel, and is working towards a solution. We will share additional updates as they become available.
Aug 1, 21:29 UTC
Investigating - Our engineering team is investigating issues related to the Cloud Control Panel. During this time, users may receive error messages when attempting to access or perform actions within the Cloud Control Panel. We apologize for the inconvenience and will share an update once we have more information.
Aug 1, 20:43 UTC
Aug 1, 2018
Resolved - Our engineering team has resolved the issue with enabling IPv6. Enabling IPv6 on an existing droplet should now be operating normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 1, 19:59 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with enabling IPv6 and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 1, 19:30 UTC
Investigating - Our engineering team is investigating an issue with enabling IPv6. During this time, you may experience the inability to enable IPv6 on an existing Droplet. We apologize for the inconvenience and will share an update once we have more information.
Aug 1, 19:03 UTC
Resolved - Our engineering team has resolved the issue with Monitoring for Droplets in our SFO2 region. Droplet Monitoring should now be operating normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
Aug 1, 17:10 UTC
Monitoring - Our engineering team has implemented a fix to resolve the issue with Monitoring for Droplets in our SFO2 region and is monitoring the situation. We will post an update as soon as the issue is fully resolved.
Aug 1, 16:41 UTC
Identified - Our engineering team has identified the cause of the issue with Monitoring for Droplets in our SFO2 region and is actively working on a fix. We will post an update as soon as additional information is available.
Aug 1, 16:14 UTC
Investigating - Our engineering team is investigating an issue with Monitoring impacting graphs for Droplets in our SFO2 region. During this time, you may not find recent monitoring data for your droplets. We apologize for the inconvenience and will share an update once we have more information.
Aug 1, 15:40 UTC